economics help

(Advanced Analysis) Given the following diagrams: Q1 = 20 bags. Q2 = 15 bags. Q3 = 27 bags. The market equilibrium price point b is $45 per bag.

The price at point a is $85 and the price at point c is $5 per bag. The price at point d is $55 and the price at point e is $35 per bag. The price at point f is $59and the price at point g is $31 per bag.

Apply the formula for the area of a triangle (Area = ½ × Base × Height) to answer the following questions.


a. What is the dollar value of the total surplus (producer surplus plus consumer surplus) when the allocatively efficient output Q1 level is being produced?

$

How large is the dollar value of the consumer surplus at that output Q1 level?
Instructions: Round your answer to one decimal place.

$

b. What is the dollar value of the deadweight loss when output level Q2 is being produced?
Instructions: Round your answer to one decimal place.

$

What is the total surplus when output level Q2 is being produced?
Instructions: Round your answer to one decimal place.

$

c. What is the dollar value of the deadweight loss when output level Q3 is produced?

$

What is the dollar value of the total surplus when output level Q3 is produced?

$
Answer #1

a. Total surplus is the area bounded by points a, b, and c. To calculate total surplus, we use the following formula for the area of a triangle: Area = ½ × Base × Height. The area between the demand curve and the supply curve for the quantity ranging from 0 to 20 is the total economic surplus. This is a triangle with a base (best read of the price axis) of $80, which is the price difference at Q = 0, or between points a and c, and a height of 20 (the number of units purchased in equilibrium). Using these values, we have a total surplus of (1/2) × $80 × 20 = $800.

The consumer surplus is the area between the demand curve and the equilibrium price line. Here we have a base of $40 (the price difference between the demand schedule price at Q = 0, which is $85, and the equilibrium price of $45). The height of the triangle is once again 20 (the number of units purchased in equilibrium). Using these values, we have a consumer surplus of (1/2) × 40 × 20 = $400.

b. Deadweight loss is the difference in total surplus between an efficient level of output Q1 and a reduced level of output at Q2. We can calculate this as the area of a triangle bounded by points bde. The base of this triangle is the difference in prices at points d and e, or $55 – $35 = $20. The height of this triangle is given by the difference in the restricted level of output of Q2 = 15 and the efficient level of output Q1 = 20, or 5 units. Thus, the area of this triangle (the deadweight loss) is equal to (1/2) × $20 × 5 = $50. The remaining total surplus can be found by subtracting the deadweight loss from the original (efficient) total surplus. This is $800 (maximum total surplus) – $50 (deadweight loss) = $750.

c. The deadweight loss from overproduction is the difference in total surplus between an efficient level of output Q1 and an additional level of output at Q3. We can calculate this as the area of a triangle bounded by points bfg. The base of this triangle is the difference in prices at points f and g, or $59 – $31 = $28. The height of this triangle is given by the difference in the additional level of output Q3 = 27 and the efficient level of output Q1 = 20, or 7 units. Thus, the area of this triangle (the deadweight loss) is equal to (1/2) × $28 × 7 = $98. The remaining total surplus can be found by subtracting the deadweight loss from the original total surplus. This is $800 (maximum total surplus) – $98 (deadweight loss) = $702. Note here that we maximize total (producer + consumer) surplus by producing the equilibrium quantity, but we lose surplus from overproduction (inefficient use of resources).

answered by: bizfan